# Multi-modal Retrieval Augmented Multi-modal Generation: A Benchmark, Evaluate Metrics and Strong Baselines

### v1

@@ Abstract
This paper investigates an intriguing task of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG). This task requires foundation models to browse multi-modal web pages, with mixed text and images, and generate multi-modal responses for solving user queries, which exhibits better information density and readability. Given the early researching stage of M2RAG task, there is a lack of systematic studies and analysis. To fill this gap, we construct a benchmark for M2RAG task, equipped with a suite of text-modal metrics and multi-modal metrics to analyze the capabilities of existing foundation models. Besides, we also propose several effective methods for foundation models to accomplish this task, based on the comprehensive evaluation results on our benchmark. Extensive experimental results reveal several intriguing phenomena worth further research.
@@
A Benchmark, Evaluate Metrics and Strong Baselines
Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
Figure 1: A typical comparison between naive RAG and our proposed M2RAG. GPT-4o is used in this case.
Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the dataset construction pipeline and Step 4 demonstrates the generation process.
@@ 1 Introduction
Retrieval Augmented Generation (RAG) has been emerged as a promising technique to enhance the factual correctness of model-generated responses by grounding on facts in external knowledge databases Lan et al. (2023); Sun et al. (2024); Li et al. (2022). Except for textual knowledge, multi-modal data, like images and videos, also contain diverse information, and it is natural to leverage these knowledge to answer user queries more accurately, leading to a concept named Multi-modal RAG or MuRAG Chen et al. (2022); Joshi et al. (2024); Ye et al. (2024). Numerous works Riedler and Langer (2024) have demonstrated that RAG with multi-modal inputs outperforms single-modal RAG. This paper studies MuRAG, but we primarily focus on how to use foundation models to generate multi-modal responses for solving user queries, which we named M2RAG task, given multi-modal inputs with mixed layout of text and images. Numerous previous works Zhu et al. (2024); Wu et al. (2023b); Li et al. (2024) have proven that multi-modal outputs effectively improve the information density and readability for users, embodying the principle—A single image is worth a thousand words. For example, Figure 1 compares plain text and multi-modal responses to the same user query about folding a paper airplane. While both responses provide similar textual explanations, the inclusion of instructional images in the multi-modal response significantly enhances readability. As a result, human readers can understand the instructions more quickly and efficiently, enabling them to successfully create a paper airplane with less effort.
Despite the emerging interest in the M2RAG task, there is a lack of systematic analysis and evaluation, given its early stage of research development. To address this gap, we propose a benchmark designed to evaluate the capabilities of existing generative models, including Large Language Models (LLMs) and Multi-modal Large Language Models (MLLMs), in completing the M2RAG task. This benchmark evaluates the generated content from both text-modal and multi-modal perspectives, incorporating four metrics for each dimension, such as fluency, relevance, image coherence, and helpfulness. The benchmark dataset is constructed through a multi-step workflow, as illustrated in Steps 1-3 of Figure 2: (1) Query Collection: We gather a diverse set of user queries spanning ten distinct topics. For each query, the top-$k$ relevant multi-modal web pages are retrieved using the Google Search API; (2) Data Cleaning: A rigorous data cleaning pipeline is applied to these web pages to extract and filter high-quality text and images, which are then used to build a reliable multi-modal database for retrieval; (3) Retrieval: A suite of retrieval methods is developed to evaluate the relevance between multi-modal data elements and the user query. Each element is assigned a relevance score to guide its selection for further processing. Additionally, we propose two approaches for tackling this task, based on comprehensive studies using the constructed benchmark: (1) Single-stage Approach: This method generates a multi-modal response directly in a single turn, with images represented as placeholders within the response; (2) Multi-stage Approach: This method first generates a text response and then interleaves relevant images, finally refines the text response with retrieved images to ensure a more cohesive and enriched output.
Extensive experimental results uncover numerous intriguing phenomena, providing valuable insights for future researches. For example, the multi-stage approach integrates more relevant images into responses, significantly outperforming the single-stage approach. Moreover, existing MLLMs demonstrate notable limitations in handling M2RAG tasks, performing 8% and 13% worse than LLMs under the single-stage and multi-stage approaches, respectively, which underscores substantial room for improvement. Importantly, compared to separate modeling approaches employed in prior works such as MuRAR Zhu et al. (2024), jointly modeling the relationships and interactions between images and text results in noticeable performance enhancements for the M2RAG task. These findings highlight the necessity of advancing joint modeling techniques to unlock the full potential of multi-modal generative models. Overall, our contributions are summarized as follows:
  • Task Formulation and Benchmark: We introduce the M2RAG task and establish a comprehensive benchmark, alongside a suite of fine-grained text-modal and multi-modal metrics to evaluate the capabilities of existing models.
  • Proposed Baselines: We propose several strong baseline methods for solving the M2RAG task, including single-stage and multi-stage approaches.
  • Experimental Insights: Our experiments demonstrate that current LLMs outperform MLLMs significantly in M2RAG tasks. Additionally, decomposing this complex task into simpler sub-tasks further enhances performance. We also observe that relying solely on images in this task can result in poor performance for certain topics, suggesting that incorporating additional modalities could yield further improvements.
@@ 2 Related Work
2.1 Text-modal Retrieval Augmented Generation (RAG)
@@
RAG Li et al. (2022) has been widely used to improve the generation quality of language models by incorporating external knowledge Lan et al. (2023), which is usually retrieved by using BM25, dense retrieval model or search engine Lee et al. (2021); Asai et al. (2023). Recently, given the increasing capabilities of LLMs to effectively process any text inputs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al. (2024).
2.2 Multi-modal RAG
@@
Although typical RAG framework demonstrates great potential, it can only handle textual inputs and cannot utilize the rich information presented in multi-modal data such as images and videos Su et al. (2023). While earlier efforts to incorporate multi-modal inputs for text generation [^1] Most existing multi-modal works mainly focus on the visual modality. often relied on carefully designed frameworks Xiao et al. (2024), a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently Chen et al. (2024); Wang et al. (2024a); Yao et al. (2024); Shen et al. (2023); Riedler and Langer (2024), with representative models including GPT-4o Hurst et al. (2024) and Llama-3.2-Vision Dubey et al. (2024). While these multi-modal models greatly expand the range of applications for LLMs, they remain limited to text-only outputs, restricting their ability to directly deliver rich, multi-modal information to users Zhu et al. (2024).
2.3 Multi-modal Generation
@@
In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts Zhu et al. (2024). Consequently, it is crucial to develop foundation models that not only generate plain text responses to user queries but also incorporate relevant multi-modal data to enhance readability and user engagement. This approach embodies the principle—A single image is worth a thousand words, emphasizing the value of visual elements in effective communication. To the best of our knowledge, MExBERT Singh et al. (2021) took the first step in this research direction, which retrieves one image given the user query and generated response. Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT Wu et al. (2023b), TextBind Li et al. (2024), Janus Wu et al. (2024) and Emu3 Wang et al. (2024b). Unlike these works, our study focuses on M2RAG task, which aims to dynamically select the multi-modal content from multiple multi-modal inputs (text or images) to construct final multi-modal responses without generating any visual elements. To the best of our knowledge, MuRAR Zhu et al. (2024) is the most closely related work to our study. However, it falls short of modeling multi-modal content holistically. Specifically, it generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data. This approach differs significantly from the setup of our proposed M2RAG task, which also emphasizes the integration of multi-modal content in the input.
@@ 3 Task Formulation
In M2RAG task, models answer each user query $Q$ with a multi-modal response based on retrieved multi-modal web pages. Suggest that multi-modal data resource $R=\{D_{1},\cdots,D_{n}\}$ represents $n$ retrieved multi-modal documents or web pages from the Internet. Each document $D_{i}$ consists of $m$ ordered elements $\{E_{i,1},E_{i,2},\cdots,E_{i,m}\}$, where each element $E_{i,j}$ can be either a textual or visual element. Elements are preserved in their original order within a document, ensuring that images are accompanied by their corresponding textual contexts. These contexts often provide rich descriptions or additional information that enhances the understanding and relevance of the images.
With the above inputs, the M2RAG task needs several steps to get the multi-modal response. Specifically, the following procedures are essential: (1) In-Doc Retrieval, (2) Generation and (3) Multi-modal Element Incorporation.
@@
While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time increases significantly as input sequences grow longer Sun et al. (2024). This challenge is more severe in the M2RAG task, where models must process both textual content and corresponding images simultaneously.
To mitigate this issue, we first propose the In-Doc Retrieval process to collect relevant elements from $R$, which are used for next step multi-modal generation.
As illustrated in Step 3 of Figure 2, we employ a retrieval model, denoted as $M_{R}$, to assess the relevance of each element $E_{i,j}$ in the dataset $R$ with respect to the user query $Q$, denoted as $M_{R}(Q,R)$. The top-$k$ most relevant elements are selected to form a refined multi-modal data resource, $R_{\text{In-Doc}}$. The refined resources serve as the inputs for generative models. Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input data for the generation process.
@@
The refined multi-modal data resource $R_{\text{In-Doc}}$ is then processed by the generative model $M_{G}$ alongside the user query $Q$ to generate the final multi-modal response:
$r=M_{G}(Q,R_{\text{In-Doc}})$ (1)
@@
Here, $r$ represents the final answer to the user query. Notably, if no visual elements are in $R_{\text{In-Doc}}$, the task defaults to the traditional RAG, since no multi-modal elements are utilized and generated.
Multi-modal Element Incorporation
Different from previous works that retrieve multi-modal data after text generation Zhu et al. (2024), we proposed a novel approach that can incorporate multi-modal data with text in a single-turn conversation. Specifically, images in the input and output are presented as placeholders. Each placeholder follows the following format: ![IMAGE_CONTENT](PSEUDO_URL). The pseudo URL identifies which image the model intends to include Wu et al. (2023a) in the output. For input images, the content is represented as tokenized image embeddings for MLLMs or as detailed image captions for LLMs. For output images, the image content is a concise caption based on the input image content which aids in the Chain-of-Thought (CoT) process Wei et al. (2022), assisting the model in understanding the image content or caption and placing it appropriately.
4 Dataset Construction
As shown in Step 1-3 of Figure 2, we elaborate three detailed steps to construct the benchmark dataset for evaluating the capabilities of foundation models in M2RAG task: (1) Query Collection, (2) Data Preparation and (3) Element Evaluation.
4.1 Query Collection
@@
The queries are sourced from the ELI5 Fan et al. (2019) dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2. Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for providing comprehensive answers. After filtering, the remaining queries are classified to ensure a balanced dataset and lay the groundwork for further studies on the topic sensitivity of the proposed methods [^2] Please refer to Appendix A for more details.. The classification process results in 11 distinct categories, as detailed in Table 1. 20 queries are sampled in each category (other than “Others”), resulting in a final benchmark dataset containing 200 queries.
ID Category Name Num. Samples
1 Society & Culture 14182
2 Science & Mathematics 14449
3 Health 10931
4 Education & Reference 776
5 Computers & Internet 6999
6 Sports 1974
7 Business & Finance 3821
8 Entertainment & Music 1946
9 Family & Relationships 820
10 Politics & Government 4638
11 Others 451
Table 1: Topics and volumes of the queries.
4.2 Data Preparation
All multi-modal data used for retrieval and generation are prepared during this step. Specifically, data for each query are processed through the following procedures: (1) Prepare Web Pages and Auxiliary Images, (2) Text Processing and (3) Image Processing.
4.2.1 Prepare Web Pages and Auxiliary Images
@@
For each collected query, we first employ Google Search API [^3] https://developers.google.com/custom-search/v1 to get the URLs of web pages given the query, and then retrieve each web page containing a mix of text and images in markdown format using JINA AI Reader [^4] https://jina.ai/reader/. Our preliminary study reveals that the images within these web pages are sometimes inaccessible, low quality, or insufficient for effective multi-modal generation. Therefore, we supplement the dataset with auxiliary images sourced from Google Image Search, using the user queries as search input to ensure relevance and enrich the visual content.
@@
Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset, as illustrated in Step 2 of Figure 2.
4.2.2 Text Processing
@@
The raw web pages often contain a significant amount of irrelevant or redundant text, like the advertisements and web links, and it is essential to remove these data and produce concise and relevant content for generation. The text processing process is divided in to three steps: (1) Embedded Image Handling: image URLs are extracted from the web pages for image collection, and placeholders are inserted in their place within the text; (2) Text Cleaning: text snippets matching pre-defined patterns, such as web links, are removed to ensure cleaner and more relevant content; (3) Text Splitting: the cleaned text content is segmented into smaller pieces using the LangChain Chase (2022) text splitter, facilitating further selection and processing.
4.2.3 Image Processing
@@
We conduct following steps to process images, ensuring they are of high quality and relevant to the user query: (1) Caching & Conversion: we download all the images with their URLs, and convert them into widely accepted formats such as JPG, PNG, GIF, or WEBP. Images that fail to be downloaded or converted are discarded; (2) Filtering: images that fall below a certain size threshold or receive a low similarity score with CLIP Radford et al. (2021) are removed. Such images are typically insufficiently informative, often consisting of icons, banners, or other non-representative visuals. (3) Deduplication: PHash algorithm-based deduplication Zauner (2010) is applied to remove duplicated or similar images, preventing redundant images for generation.
4.3 Element Evaluation

### v2

# Multi-modal Retrieval Augmented Multi-modal Generation: Datasets, Evaluation Metrics and Strong Baseliness

@@ Abstract
We present a systematic investigation of Multi-modal Retrieval Augmented Multi-modal Generation (M2RAG), a novel task that enables foundation models to process multi-modal web content and generate multi-modal responses, which exhibits better information density and readability. Despite its potential impact, M2RAG remains understudied, lacking comprehensive analysis and high-quality data resources. To address this gap, we establish a comprehensive benchmark through a rigorous data curation pipeline, and employ text-modal metrics and multi-modal metrics based on foundation models for evaluation. We further propose several strategies for foundation models to process M2RAG effectively and construct a training set by filtering high-quality samples using designed metrics. Our extensive experiments demonstrate the reliability of our proposed metrics, a landscape of model performance within our designed strategies, and show that our fine-tuned 7B-8B models outperform the state-of-the-art GPT-4o model. Additionally, we perform fine-grained analyses across diverse domains and validate the effectiveness of our designs in data curation pipeline. All resources, including codes, datasets, and model weights, will be publicly released.
@@
Datasets, Evaluation Metrics and Strong Baseliness
Zi-Ao Ma1, Tian Lan1, Rong-Cheng Tu2, Yong Hu3 Yu-Shi Zhu1, Tong Zhang1, Heyan Huang1, Xian-Ling Mao1thanks:  Corresponding author 1School of Computer Science and Technology, Beijing Institute of Technology, China 2Nanyang Technological University, Singapore, 3WeChat AI, Tencent Inc., China {maziaoylwt,lantiangmftby}@gmail.com, rongcheng.tu@ntu.edu.sg rightyonghu@tencent.com, maoxl@bit.edu.cn https://github.com/maziao/M2RAG
Figure 1: A typical comparison between naive RAG (upper) and our proposed M2RAG (lower). The generative model is GPT-4o in this case.
@@ 1 Introduction
Retrieval Augmented Generation (RAG) Lan et al. (2023); Sun et al. (2024); Li et al. (2022) and its multi-modal extensions  Chen et al. (2022); Joshi et al. (2024); Ye et al. (2024) enhance foundation models by incorporating external knowledge and multi-modal data. While these methods improve response quality, they remain limited to textual outputs, which may fail to provide sufficient clarity in some scenarios, such as instructional guides or spatial reasoning. For instance, as shown in Figure 1, a text-only response explaining paper airplane folding steps may be difficult for users to follow without accompanying visual illustrations.
As the saying goes, A single image is worth a thousand words. Visual elements can significantly enhance comprehension, especially in instructional and knowledge-intensive domains where textual descriptions alone may be insufficient. Motivated by this, the Multi-modal Retrieval-Augmented Multi-modal Generation (M2RAG) has been introduced  Zhu et al. (2024); Wu et al. (2023); Li et al. (2024), a novel task that retrieves and integrates multi-modal content to produce responses with a mixed layout of text and image. By integrating relevant images into textual responses, M2RAG enhances comprehension and usability across various application scenarios. As illustrated in Figure 1, M2RAG not only generates textual explanations for step-by-step folding a paper airplane but also seamlessly incorporates illustrative images where necessary, significantly improving user understanding. Compared to existing RAG methods, M2RAG introduces new challenges in multi-modal reasoning and synthesis. It requires foundation models to achieve a deeper understanding of multi-modal data, accurately capture content and cross-modal relationships, and generate coherent responses that seamlessly integrate text and images.
However, given the early research stage, there is a lack of systematic analysis and high-quality data resources for the M2RAG task. To fill this gap, we make following contributions: (1) Benchmark: A comprehensive benchmark with 10 topics constructed by our proposed rigorous data curation pipeline; (2) Evaluation Metric: A suite of reliable text-only and multi-modal evaluation metrics based on Large Language Models (LLMs) and Multi-modal LLMs (MLLMs); (3) Generation Strategy: Two generation strategies for effectively tackling M2RAG task: single-stage and multi-stage approaches; and (4) Training Dataset: A training dataset constructed by filtering high-quality samples using our designed multi-modal evaluation metrics, which is used to improve the performance of 7B-8B LLMs and MLLMs.
Based on our extensive experiments and automatic evaluation of advanced LLMs and MLLMs, we present four key findings: (1) Metric Reliability: The used evaluation metrics exhibit strong correlations with human judgments, even outperform inner-correlation of human annotators in evaluating image helpfulness dimension. This validates the reliability of our automatic evaluation framework; (2) Model Performance Analysis: Experimental results on 12 advanced LLMs and MLLMs provides valuable insights for M2RAG task. For example, LLMs consistently outperforms MLLMs, revealing MLLM’s limited multi-modal understanding and generation abilities. Besides, multi-stage strategy integrates more relevant images into responses, exhibiting better overall quality; (3) Effectiveness of Training Dataset: Fine-tuning 7B-8B LLMs and MLLMs on our curated dataset yields performance that exceeds GPT-4o, underscoring the quality and effectiveness of our training dataset; and (4) Data Curation Benefits: Ablation studies quantify cross-domain performance variations and verify the positive contributions of our designs in benchmark data curation pipeline, including detailed context for images and the inclusion of the auxiliary images. These observations and phenomena promote an in-depth understanding of M2RAG. We hope our data resources and these discoveries could spur future research in this field.
@@ 2 Related Work
Text-modal RAG
@@
RAG Li et al. (2022) has been widely used to improve the generation quality of language models by incorporating external knowledge Lan et al. (2023), which is usually retrieved by using BM25, dense retrieval model or search engine Lee et al. (2021); Asai et al. (2023). Recently, given the increasing capabilities of LLMs, there is an emerging way to simply concatenate all retrieved data into the context of LLMs for generation Sun et al. (2024).
Multi-modal RAG
@@
Despite the great potential of RAG, it can only handle textual inputs and cannot utilize the rich information in multi-modal data such as images and videos Su et al. (2023). While earlier efforts to incorporate multi-modal inputs for text generation often relied on carefully designed frameworks Xiao et al. (2024), a new paradigm has emerged that emphasizes pre-training MLLMs to accomplish this task more directly and efficiently Chen et al. (2024); Wang et al. (2024a); Yao et al. (2024); Shen et al. (2023); Riedler and Langer (2024), with representative models including GPT-4o Hurst et al. (2024) and Llama-3.2-Vision Dubey et al. (2024). While these multi-modal models greatly expand the range of LLMs’ applications, they remain limited to textual outputs, restricting their ability to deliver rich multi-modal responses to users Zhu et al. (2024).
Multi-modal Generation
@@
In real-world scenarios, humans naturally interact with multi-modal data, such as browsing web pages that combine text, images, and videos in mixed layouts Zhu et al. (2024). Consequently, it is crucial to develop foundation models that not only generate plain text responses to user queries but also incorporate relevant multi-modal data to enhance readability and user engagement. This approach embodies the principle—A single image is worth a thousand words, emphasizing the value of visual elements in effective communication. To the best of our knowledge, MExBERT Singh et al. (2021) took the first step in this research direction, which retrieves one image given the user query and generated response. Beyond retrieving the images, recent works also focus on generating the corresponding images for model-generated responses, like Next-GPT Wu et al. (2023), TextBind Li et al. (2024), Janus Wu et al. (2024) and Emu3 Wang et al. (2024b). Unlike these works, our study focuses on M2RAG task, which aims to dynamically select the multi-modal content from multiple multi-modal inputs (text or images) to construct final multi-modal responses without generating any visual elements. To the best of our knowledge, MuRAR Zhu et al. (2024) is the most closely related works to our study. However, they falls short of modeling multi-modal content holistically. For example, MuRAR generates initial responses as plain text based solely on user queries and retrieved textual content, without incorporating the corresponding multi-modal data. These approaches differs significantly from the setup of our proposed M2RAG task, which emphasizes the understanding the content and relationship among multi-modal input data.
@@ 3 Task Formulation
In M2RAG task, models generate a multi-modal response $r$ to each user query $Q$ by summarizing a retrieved multi-modal knowledge base $K=\{D_{1},\cdots,D_{n}\}$, consisting of $n$ multi-modal documents or web pages from the Internet. Each document $D_{i}$ consists of $m$ ordered elements $\{E_{i,1},E_{i,2},\cdots,E_{i,m}\}$, where each element $E_{i,j}$ can be either a text paragraph or an image. The original order of elements within a document is preserved to ensure images remain paired with their associated textual context, providing rich descriptions that enhance image comprehension. The process of generating the multi-modal response involves two key steps: In-Doc Retrieval and Generation.
@@
While previous studies have demonstrated the strong capabilities of LLMs to answer user queries grounded in retrieved documents, their inference time grows significantly with longer input sequences Sun et al. (2024). This challenge is more severe in the M2RAG task, where models must handle extensive textual and visual data across multiple documents. To address this challenge, we introduce the In-Doc Retrieval, which selects the most relevant and useful elements from $K$ to reduce inference costs. As illustrated in Figure 2 (Step 3), a retrieval model $M_{R}$ assesses the relevance of each element $E_{i,j}$ in the knowledge base $K$ with respect to the user query $Q$: $M_{R}(Q,K)$. The top-$k$ most relevant elements are selected to form a refined and concise knowledge base, $K_{\text{In-Doc}}$. Importantly, when a visual element is selected, its associated textual context is also retrieved to ensure coherence and provide richer input information for the generation process. The implementation details of the retrieval model is provided in Section 4.1.3.
@@
The refined knowledge base $K_{\text{In-Doc}}$ is then processed by the generative model $M_{G}$ alongside the user query $Q$ to generate the final multi-modal response:
$r=M_{G}(Q,K_{\text{In-Doc}})$ (1)
@@
Here, $r$ represents the final answer to the user query. Notably, if no visual elements are in $K_{\text{In-Doc}}$, the task defaults to the traditional RAG, since no multi-modal elements are utilized and generated.
Figure 2: The overall framework of our proposed dataset construction and M2RAG pipeline. Step 1-3 represent the benchmark data curation pipeline and Step 4 demonstrates our proposed generation strategies.
4 Datasets and Methodology for M2RAG
We describe details of our work: (1) Benchmark (Section 4.1); (2) Evaluation Metrics (Section 4.2); (3) Generation Strategy (Section 4.3); and (4) Training Dataset (Section 4.4).
4.1 Benchmark Construction
Our benchmark are established by running following steps: (1) Query Collection; (2) Data Preparation; and (3) Retrieval.
4.1.1 Query Collection
@@
The queries are sourced from the ELI5 Fan et al. (2019) dataset and undergo a series of filtering and classification procedures, as outlined in Step 1 of Figure 2. Specifically, we first exclude queries that do not require rich references and then discard those that do not necessitate visual information for answers. The filtered queries are then classified into 11 categories to ensure dataset balance and support further investigation of the proposed methods’ topic sensitivity. [^1] See Appendix B for details.
4.1.2 Data Preparation
Subsequently, we crawl the multi-modal data related to each query and post-process them by conducting following steps (Step 2 of Figure 2): (1) Web Page and Auxiliary Image Crawling; (2) Text Processing; and (3) Image Processing.
Web Page and Auxiliary Image Crawling
@@ 4.1.2 Data Preparation
For each collected query, we first employ Google Custom Search API [^2] https://developers.google.com/custom-search to get the URLs of web pages given the query, and then retrieve each web page containing a mix of text and images in markdown format using JINA AI Reader [^3] https://jina.ai/reader/. Our preliminary study reveals that the images within these web pages are sometimes inaccessible, of low quality, or insufficient for effective multi-modal generation. Therefore, we supplement the dataset with auxiliary images sourced from Google Image Search, using the user queries as search input to ensure relevance and enrich the visual content.
@@
Since the qualities of collected web pages and images are not guaranteed, we further propose Text Processing and Image Processing to post-process these multi-modal data to improve the quality of our dataset.
Text Processing
@@
Raw web pages often include irrelevant or redundant text, such as advertisements and web links, which must be removed to produce concise and relevant content for generation. The text processing pipeline consists of three steps: (1) Embedded Image Handling: Extract image URLs from web pages for image collection and replace them with placeholders in the text; (2) Text Cleaning: Remove text snippets matching predefined patterns, such as web links, to ensure cleaner and more relevant content; (3) Text Segmentation: Employ the text splitter Chase (2022) to split the cleaned text into smaller pieces, enabling efficient selection and processing.
Image Processing
@@
We conduct following steps to process images to ensure they are of high quality and relevant to the user query: (1) Caching & Conversion: All images are downloaded using their URLs and converted into widely accepted formats such as JPG, PNG, GIF, or WEBP. Images that cannot be successfully downloaded or converted are discarded; (2) Filtering: images smaller than a certain threshold or with a low CLIP-based Radford et al. (2021) similarity score to the query text are removed. Such images often consist of non-representative visual contents, such as icons, banners, etc. (3) Deduplication: duplicate or highly similar images are removed using PHash Zauner (2010) algorithm.
4.1.3 Element Evaluation